OpenStudy (anonymous):

A table is on sale for 65% off. The sale price is $308 . What is the regular price?

OpenStudy (anonymous):

Say the regular price is x. 65 % off the price is then x*0.65, which equals $308. Now in math: x*0.65 = 308 so, x*0.65/0.65 = 308/0.65 => x = 308/0.65

OpenStudy (anonymous):

Was that understandable?

OpenStudy (anonymous):

No i am still confused on how to get the regular price

OpenStudy (anonymous):

1 % means 1 hundredth 65 % means 65 hundrendths. Here we have 65 hundrenths of the regular price equaling 308 dollars, which is the sales price. If we divide 308 dollars by 0.65 we get the regular price.

OpenStudy (anonymous):

The regular price would be higher than the sales prices, naturally

OpenStudy (anonymous):

Anything unclear??

OpenStudy (anonymous):

308/0.35 = 880

OpenStudy (anonymous):

"65% off".. yeaaaah

OpenStudy (anonymous):

Sorry, I misunderstood.

OpenStudy (anonymous):

The sale price is 65% less than the regular price. So, the sale price is only 100%-65%=35% of the regular price. Sale price =35% Regular price The sale price is $308 , and so we have the following. 308 =35%x Regular price 308 =35%= Regular price To find the regular price, we divide 308 by 0.35 . Regular price=305/0.35=880

OpenStudy (anonymous):

@saraward070911 Thank you for the medal.

OpenStudy (anonymous):

I interpreted the question as $308 being 65 %, not being 65 % off

OpenStudy (anonymous):

Your Welcome @Robtobey